Two planes leave an airport at the same time. Their speeds are 130 mph and 150 mph, and the angle between their courses is 36 degrees. How far apart are they after 1.5 hours?
Answer by Alan3354(69443)

Two planes leave an airport at the same time. Their speeds are 130 mph and 150 mph, and the angle between their courses is 36 degrees. How far apart are they after 1.5 hours?
--------------------
The distances they've flown are 2 sides of a triangle.
a = 130*1.5 = 195 miles
b = 150*1.5 = 225 miles
The angle between the sides is 36º
Use the Cosine Law to find the 3rd side:
c^2 = a^2 + b^2 - 2*a*b*cos(36)